a number consists of two digits the digits in the ten's place is 3 times the digit in the unit place if 54 is subtracted from the number the digits are reversed . The number is
Answer by ankor@dixie-net.com(15746)

a number consists of two digits the digits in the ten's place is 3 times the digit in the unit place if 54 is subtracted from the number the digits are reversed.
The number is
:
Let the two digit number = 10x + y
:
"the ten's place is 3 times the digit in the unit place"
x = 3y
:
"if 54 is subtracted from the number the digits are reversed."
10x + y - 54 = 10y + x
10x - x = 10y - y + 54
9x = 9y + 54
simplify, divide by 9
x = y + 6
Replace x with 3y
3y = y + 6
3y - y = 6
2y = 6
y = 3
then
x = 3(3)
x = 9
:
93 is the number
:
:
Check solution in the statement:
"if 54 is subtracted from the number the digits are reversed."
93 - 54 = 39